what line is perpedincular to the line y=2/3x+3?
please help me solve this question or show me the process of finding an answer!
thank you

Answer by Alan3354(69443)

what line is perpedincular to the line y=2/3x+3?
-----------
The slope of the line y=2/3x+3 is 2/3.
All lines perpendicular to it have a slope of -3/2, the negative reciprocal.
-------------
y = (-3/2)x + b is perpendicular, for b = any value.